OpenStudy (anonymous):

Is there a formula for the product of a sequence like how there is for the sum of an arithmetic and geometric series?

OpenStudy (anonymous):

a product of a sequence is usually a factorial

OpenStudy (anonymous):

Example: \[\prod_{i = 3}^{7}2^{i}\]

OpenStudy (lgbasallote):

\[n! = n \times (n-1) \times (n-2) \times \cdots 1?\]

OpenStudy (anonymous):

okay in this case you can take the natural log of both sides to make that a summation

OpenStudy (anonymous):